A car is to be hoisted by elevator to the third floor of a parking garage, which is 36 ft above the ground. PartA If the elevator can accelerate at 0.7 ft/s2, decelerate at 04 ft/s2, and reach a maximum speed of 8 ft/s, determine the shortest time to make the lift, starting from rest and ending at rest.
Answer #1

PEvdZAC CH) 29--0sy iuy 29-8129 2-Co7)24 Vmax= 579 HAくきHk 5-796 0+ o 1tj = 5.796-o-utz ヒュ= 14491 s 6etiat 22-775